1) Antiderivative of -4csc(x)cot(x)dx :Show all work

 

2) Antiderivative of -1cot^2(x)dx :Show all work

 

3) Antiderivative of ((6/x^3)+cube root of x^2)dx :Show all work

 

4) Write the Antiderivative and definite integral (leave in terms of e if applicable) --> the integral from -1 to 3 (10e^x-4x)dx

 

5) Write the Antiderivative and definite integral(leave in radical form if applicable)--> the integral from (pi/2) to (5pi/6) (-3cos(x)-4sin(x))dx




1
Expert's answer
2021-03-03T10:47:05-0500

1) "\u222b-4csc(x)cot(x)dx"

"-4(∫\frac{1}{sin\left(x\right)}\frac{\cos{\left(x\right)}}{\sin{\left(x\right)}}\ )\ dx"

"-4∫\frac{\cos{\left(x\right)}}{\sin^2{(x)}}\ dx"

let u = sin(x)

du = cos(x) dx

"-4∫\frac{du}{u^2}"

"\frac{4}{u}+c"

"\frac{4}{\sin{\left(x\right)}}+c\"

"4\csc{\left(x\right)}+c"


2) "∫-1\cot^2{\ \left(x\right)\ dx\ }"

"-1∫\frac{\cos^2{\left(x\right)}}{\sin^2{\left(x\right)}}\ dx"

"-1∫\ \frac{{1-\sin}^2{\left(x\right)}}{\sin^2{\left(x\right)}}dx"

"-1∫\ \frac{1}{\sin^2{\left(x\right)}}dx\ +\ 1\int dx"

let t = tan(x)

"dt = \ sec^2\left(x\right)\ dx"

"sin(x) = \frac{t}{\sqrt{1+t^2}}"

"\frac{1}{\sin^2{\left(x\right)}}=\frac{1+t^2}{t^2}"

"\cos{\left(x\right)}=\frac{1}{\sqrt{1+t^2}}"

"\sec^2{\left(x\right)}=1+t^2"

"dx=\frac{dt}{1+t^2}"

"-1∫\ \frac{1}{\sin^2{\left(x\right)}}dx+1\int dx=\ -1\int\frac{1+t^2}{t^2}\frac{dt}{1+t^2}+\int\frac{dt}{1+t^2}"

"-1∫\frac{dt}{t^2}\ +\tan^{-1}{\ }(t) + c2"

"\frac{1}{t}+ \tan^{-1}{\ }\left(t\right) + c"

"\frac{1}{\tan{\left(x\right)}}+\tan^{-1}{\ }\left(\tan{\left(x\right)}\right)+c"

"\cot{\left(x\right)}+x+c"


3) "∫(6x^{-3}+x^\frac{2}{3})\ dx\"

"-3x^{-2}+\frac{3x^\frac{5}{3}}{5} + c"


4) "\int_{-1}^{3}{(10}e^x-4x\ )\ dx"

"(10e^x-2x^2)"

substituting the limits

"\left(10e^3-2\left(3^2\right)\right)-(10e^{-1}-2\left(-1\right)^2)"

"10\left(e^3-e^{-1}\right)-16"


5) "\int_{\frac{\pi}{2}}^{\frac{5\pi}{6}}\left(-3\cos{\left(x\right)}-4\sin{\left(x\right)}\ \right)dx"

"(-3\sin{\left(x\right)}+4\cos(x))"

substituting the limits,

"(-3\sin{\left(\frac{5\pi}{6}\right)}+4\cos{\left(\frac{5\pi}{6}\right)}\ )-(-3\sin{\left(\frac{\pi}{2}\right)}+4\cos(\frac{\pi}{2}))"

"-3\left(\frac{1}{2}\right)+4\left(-\frac{\sqrt3}{2}\right)+3(1)-4(0)"

"\frac{3-4\sqrt3}{2}"